ES&H    496 days w/o lost workday
             (ballpark calc: 500 days * 200 staff * 10 hrs/day = 1,000,000
hours)

CDF     working on power info request
            Many workers on online rate problem to final storage. Not clear
            who has lead. Expect it somewhere in CDF.

CMS     working procurement plans. Power estimates w/in 2 weeks.

D0      Last week farm yield was lower than expected. Why? Investigating
                luminosity ? p14 going in and should have better
                        performance for high lum events.
        50% utilization of CAB with significant med queue load. Suspect
                it is spot MC use. Likely to be conference driven and
                to ramp up soon.
        Expect large use of pick events due to various problems (code,
                data, ...) Will stress system, particularly the db
                servers. Expect stress to be driven by lepton/photon.

EAG     Windows Server 2003 install went well. Database guy here this week
        Working plans to move file servers to FCC.

ES      Minos instructed on how to use dCache. May see burst of activity.
        Minos specific FNALU batch instructions also newly made. May see
                increase of use.

CCF     17TB tape movement day sets new record
        10TB/day sustained rate in dCache pilot also good news
        Large SAM volume seen in CDF shows that ramping up (2TB day
                record)

        At least 1 5 minute interval of 300+ Mbps offsite network usage
                (50% of link ?) Raises question of steady to peak specs.

        Working updates to CDF drives. No new issues.

        D0 fiber to desktop plans being developed.
        Most material for CDF conference rooms have arrived. Beginning
                construction.
        Blocked on miniBoone WH control room install pending current conf.
                room gear removal.

        NetBIOS attacks have ramped up as expected. 2 compromises of
                unblocked machines. Followup with inspections of the rest.

CSS

CSS OPS Report 2003-05-19

PC Mangers meeting - Wednesday, May 28th, 9-10am, FCC2A/B

Operations meeting with SGI - Tuesday, May 20, 9:30am, FCC2A

_______________________________________________________________________

1: ES&H - Environmental, Safety & Health

1.1 APC power strips

Bruce Karrels is inspecting and repairing (as necessary) the APC power
strips used in some of the Atipa Farms racks. He has found two line cord
connection assemblies that have overheated (d0cs017-176).


2: ELS - Equipment Logistics Services

2.1 SGI O2000 hw/sw maintenance contract renewal

The requisitions for the renewal are in the MISER.  With the
decommissioning of part of d0mino (128cpus & 1 meta router), the
division has reduced the maintenance charges ~ $70k off what was
forecasted.

3: ESS - Electronics Support Services

3.1 Beams - Beam Position Monitor

209 Pre-Amps are modified, excluding installation of daughter cards
(calibration board).  66 additional pre-amps will need to be modified to
meet the total quantity required of 275. 50 EchoTek boards received
today for initial testing by ESS.


4: DSG - Database Systems Group

4.1 CDF oracle machines

N. Stanfield - cdfrep01 (fcdflnx1) was upgraded to 9.2.0.3 per last
security alert.  cdfonprd (b0dau36) is scheduled for upgrade on Tuesday
May 20. This is the last database that needs the patch. The OEM (oracle
enterprise manager) pc's were also upgraded to 9203 as well.

N. Stanfield rebuilt a new composite index on cdfonprd (b0dau35) that
interrupted Icicle but it solved their query problem.

4.2 MMS interface

S. Jones fixed a bug in the mmsInterface.pl code that was sticking
carriage returns and spaces into requisition lines. This prevented a
purchase order from being created on the mms side.

5: SCS - Scientific Computing Support

5.1 D0 Farms

SCS is seeing an increase of disk errors on new D0 farms after upgrade.
They think this because they changed the partitions on data disk and are
using areas not previously used. They are actively scanning these areas
for bad block errors, but marginal ones are not being detected by the
diagnostic activity. They are investigating better methods of testing
for errors.

5.1 SUN Technology Meeting

        Sun has asked to come out and update us on new technology. They
want to present a roadmap and also talk about throughput computing. The
presentation will be held in fcc2b on the Wednesday May 21st at 1:30.
Mail Stan for details or questions

6: CSI - Core Server and Infrastructure

6.1 Newservers

Info4 and info1 had problems replicating newsgroups Monday and Tuesday.
Rebooting BOTH servers fixed the problem.

6.2 Mail Lists

Robert Kennedy attempted to post the same message to many CDF lists. His
message was marked as SPAM by Listserv and he was added to the LISTSERV
internal restricted sender group. A workaround was used so Robert could
continue posting.

7: TOC - Technical and Office Computing Support

7.1 Computer Security

TOC checked their exempted servers for files that would have confirmed a
similar break-in to the Beams Div server.  No files found and Computer
Security advised.

TOC was also advised of unusual offsite traffic to CSDDEV over last
weekend. The system had been hacked.  The system was taken offline and
Computer Security advised.  The system was rebuilt from scratch and all
SP and patches applied and now will be put into the Fermi W2K domain.

7.4 CSDserver1

The CSDserver1 upgrade to W2K is scheduled for 5-31.

Questions:
        Did we get SGI support to resolve D0 fileserver trouble.
        D0 happy. Problem not reoccured since last maint.

        Has D0 L3 been alerted to the power dist. problems seen in
        D0 offline farm ? Could have same trouble. Amber will followup.

CEPA
        Lots of dbserver activity and going well.
        looking for pilot users for disk (?) side caching.
        JIM deployment started last week and tests between
                ClueD0 and CAB are underway.

        Failure mode of farms power distribution mode understood.

Planning
        Effort reports due next Tuesday.

Projects
        Photos solicited for new CD homepage
        Zeno media working on rollover problem.
        Wed 1pm project status meeting will focus on new webpages.

Operations
        Crating of CDF ADIC underway and ahead of schedule.
        UPS upgrade request is in to DOE. Broken in to 2 years.
                Power upgrade in FY03 request.
        Phase 2 capacity planning report given last Friday. Will
                make available to interested CD parties. Ask Gerry.

 

 

-Dane Skow

 

Addendum:

The official number on the Lab Safety Committee is 660,701 hours; PPD is
leading with 1,082,344. Looks like we've got another 250 or so days to go to
break 1,000,000, by their calculation.

http://www-esh.fnal.gov/pls/default/lsc.html

-- Mark K.